# Learned Indexes with Distribution Smoothing via Virtual Points

### v1

@@ VI-C Results on Read-write Workloads
Index size. The index size overhead decreases after each batch of insertions, as shown in Fig. 12. This is because the initial gaps left by the virtual points are gradually filled up by the inserted points, hence improving the overall space utilization. Again, the index size overhead is at or below 10%, emphasizing the space efficiency of CSV.
(a) LIPP
(b) ALEX
Figure 12: Percentage of storage space increase vs. insertion batches
Query time (other distributions). The query times for the random and Zipfian distributions, as shown in Figs. 13 to 16, increase gradually with more data inserted into the index structures, while applying CSV again helps reduce the average query times.
(a) CSV
(b) Original
Figure 13: LIPP: Absolute average query time vs. insertion batches (random)
(a) CSV
(b) Original
Figure 14: ALEX: Absolute average query time vs. insertion batches: original structures (random)
(a) CSV
(b) Original
Figure 15: LIPP: Absolute average query time vs. insertion batches (Zipfian)
(a) CSV
(b) Original
Figure 16: ALEX: Absolute average query time vs. insertion batches (Zipfian)
@@
Insertion time. Fig. 17 shows the insertion times of the CSV-enhanced indexes relative to those of the original indexes. Using CSV helps improve the insertion times in most cases because the gaps left by the virtual points can be reused for insertions. CSV could also lead to higher insertion times in a few cases. This could be attributed to the fact that there are more keys at the upper levels of the CSV-enhanced indexes which may lead to more collisions with the insertions.
(a) LIPP
(b) ALEX
Figure 17: Insertion time vs. insertion batches
Robustness to insertion. Figs. 18 to 21 show the data reduced from the CSV as a percentage of the data in the original index structure for each level. It can be seen that fewer keys are inserted to lower levels (negative values) and more keys are promoted to higher levels (positive values). This suggests that the indexing functions created by CSV are more robust to the insertions as there are fewer predictions to the same position, allowing more keys to be placed in the same node. In a very limited number of instances, the CSV-enhanced indexes have more data in some lower levels. This could be attributed to the inserted data having conflicts with the existing keys. However, the maximum such extra keys added in lower levels by CSV is only 22 keys in Fig. 20(a) for level 8.
(a) LIPP
(b) ALEX
Figure 18: Data reduction percentage on Facebook by level vs. insertion batches
(a) LIPP
(b) ALEX
Figure 19: Data reduction percentage on Covid by level vs. insertion batches
(a) LIPP
(b) ALEX
Figure 20: Data reduction percentage on OSM by level vs. insertion batches
(a) LIPP
(b) ALEX
Figure 21: Data reduction percentage on Genome by level vs. insertion batches
VII Conclusion
@@
We addressed the issue of indexing data of complex distributions by modifying the key set to be more favourable for index learning, instead of developing yet another indexing function or structure. To achieve this, we proposed a CDF smoothing technique via insertion of virtual points. Further, we proposed an algorithm named CSV to utilize this technique on existing hierarchical learned index structures, to improve the query time for keys in lower levels of these index structures. The proposed algorithm is implemented on three recent learned indexes, which are evaluated on real-world datasets. The results show significant query performance improvements, i.e., up to 34%, with a controllable and low storage space overhead.

### v2

@@ 6.3. Results on Read-write Workloads
Index size. The index size overhead decreases after each batch of insertions, as shown in Figs. 10(b) and 10(e). This is because the initial gaps left by the virtual points are gradually filled up by the inserted points, hence improving the overall space utilisation. Again, the index size overhead is at or below 10%, emphasising the space efficiency of CSV. For ALEX index, the storage increase is almost negligible ($<0.5\%$). In some cases, the storage size of the CSV-enhanced ALEX is even lower than the original index, as the original ALEX index may need to create more new nodes to host the insertions which outweighs the space overhead of CSV.
@@
Insertion time. Figs. 10(c) and 10(f) show the average insertion times of the CSV-enhanced indexes relative to those of the original indexes. Using CSV helps improve the insertion times in some cases because the gaps left by the virtual points are reused for insertions. CSV could also lead to higher insertion times in other cases. This could be attributed to the fact that there are more keys at the upper levels of the CSV-enhanced indexes which may lead to more collisions with the insertions, which requires new index node creation. Overall, the insertion times of the CSV-enhanced indexes are on par to the original indexes.
7. Conclusion
@@
We addressed the issue of index learning over data of complex distributions by a CDF smoothing technique to modify the key set, instead of developing yet another indexing function or structure. We proposed an algorithm named CSV to utilize this technique on existing hierarchical learned index structures, to improve the query time for the keys in lower levels of these index structures. The proposed algorithm is implemented on three recent learned indexes, which are evaluated on real-world datasets. The results show significant query performance improvements, i.e., up to 34%, with a controllable and low storage space overhead.